Bob makes widgets. Variable costs per unit are $4. Fixed cost per unit (at an output level of 100) are $2

per unit. The normal sales price per unit is $10. A customer approaches Bob offering to buy 40 widgets

for $8 each. Assume Bob has excess capacity.

1) What is the effect on operating income if he accepts the order?

2) Assume that, to fill the special order, Bob must buy and completely use up (no future use) a special

material for $100. What is the effect on operating income if he accepts the order?

3) Assume the need for the special material as in #2. Also, assume there is no excess capacity. What

price would Bob need to charge for the special order so that operating income will be the same whether

or not he accepts the special order?

Discontinue a product line or segment decision

Beta makes product A & product B. Recent performance data is as follows

A (based on sales of 100 units)

B(based on sales of 100

units)

Company Total
Sales Revenue 800 500 1300
-VC (400) (200) (600)
=CM 400 300 700
-FC (600) (200) (800)
=Oper.Income (200) 100 (100)

4) If Beta discontinues the A product line, what will be the effect on total company operating income?

Assume all A fixed costs are avoidable if he discontinues it.

5) Assuming only two-thirds of A’s FC are avoidable, what would be the effect on operating income if A is

discontinued?

6) What would be the effect on operating income if A is discontinued, assuming only two-thirds of A’s FC

are avoidable and that sales of B would decrease by 10% if A is discontinued?
Answer #1

Only variable cost should be considered for special order. (ie. $4)

1) Operating income from special order = 40 X ($8 - $4) = $160

2) Operating income from special order = [40 X ($8 - $4)] - $100 = $60

3) To maintain same operating income from special order, the Selling price should be $10 plus Cost of special material. Cost of special material per unit = $100 / 40 = $2.5

Selling price = $10 + $2.5 = $12.5

4) If All the fixed cost of A can be avoided, The total operating income will be operating income of B. ie $100

$100 is the Beta operating income.

5) Unavoidable fixed cost equal to = 1/3 X $600 = $200

Total Operating income (loss) = $100 - $200 = ($100)

6) Percentage reduction in sales will reduce the contribution margin in the same proportion.

Contribution margin = $300 - 10% = $270

Operating income of B = $270 - $200 = $70

Unavoidable fixed cost equal to $200

Total operating income (loss) = $70 - $200 = ($130)